A physician orders KCI 40 mEq to be added to D5W 1000 mL and administered over 4 hr. The injection solution on hand is KCl 2 mEq/mL. How many milliliters of KCl should be added? a. 40 mL b. 8 mL c. 16 mL d. 20 mL ·You have been asked to prepare 1 L of a 2% (w/v) solution. You have a 100% solution in stock. How much diluent will be needed? a. 2.0 ml b. 20 mL c. 200 ml d. 980 mL
Answer #1

ANSWER.

* Calculation of milliliters of KCL added.

         Formula

           Amount of KCL added (ml) = Amount desired /Amount in hand * Quantity.

                 Amount desired(D) = 40 mEq.

             Amount in hand(H) = 2 mEq.

                      Quantity(Q) = 1 mL.

              As per formula

                  Amount of KCL is added = D / H * Q

                                                              = 40 / 2 * 1

                                                        d    = 20 mL.

* Calculation of amount of dilutent needed.

       Amount of solution present in the stock =100 %

      Amount of solution need to be prepared = 1L of 2%(w/v) solution.

                                                                       = 1000 mL of 2%(w/v).

         So

          Amount of dilutent needed = (100% - 2%) 1000 ml.

                                                         = 98% * 1000 mL.

                                                           = 98/100 * 1000 mL.

                                                           = 0.98 * 1000 mL.

                                                  d   = 980 mL.